OpenStudy (anonymous):

Two ants are at a common point at time t=0, the first ant starts crawling along a straight line at the rate of 4 ft/min. Two minutes later, the second ant starts crawling in a direction perpendicular to that of the first, at a rate of 5 ft/min. How fast is the distance between them changing when the first insect has traveled 12 feet? Thank you!

OpenStudy (some.random.cool.kid):

WELCOME TO OS!!!!

OpenStudy (anonymous):

That is not a vaild answer

OpenStudy (some.random.cool.kid):

lol

OpenStudy (some.random.cool.kid):

its a regulation most OS'ers say this when someone new comes on xD

OpenStudy (some.random.cool.kid):

hmm..

OpenStudy (some.random.cool.kid):

@sammixboo

OpenStudy (some.random.cool.kid):

@sleepyjess

OpenStudy (some.random.cool.kid):

@DanJS

OpenStudy (anonymous):

Who are they?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Let's draw a diagaram 

OpenStudy (anonymous):

We have ant 1 moving at a rate of 4 feet of every minute of time, hence the multiplication. The t+2 comes from the 2 minute headstart that ant 1 had. Using pythagoras, we can get an expression for the distance between the two ants with respect to time. \[distance = \sqrt{(4\times(t+2))^2+(5\times t)^2}\] \[D (t) = \sqrt{(4t+8)^2+(5t)^2}\] \[D(t) = \sqrt{14t^2+64t+64+25t^2}\] \[D(t) = \sqrt{39t^2+64t+64}\]

OpenStudy (anonymous):

The question asks for D'(t) after ant 1 (represented by y) has travelled 12 feet. So let's let y = 12 y = 4(t+2) 12 = 4(t+2) (12/4)-2 = t t = 1 So the question is asking for D'(1). I'm assuming you know how to derive this so: \[D'(x) = \frac{ 78t+64 }{ \sqrt{39t^2+64t+64} }\] Now we just have to substitute it in. Yay! \[D'(1) = \frac{ 78 \times 1+64 }{ \sqrt{39\times1^2+64\times1+64} }\] \[D'(1) = 10.99feet/s\] Hope this helps